Mathematics 242 – Principles of Analysis Definitions Quiz 7 – November 18, 2005

Your Name:

1. What is the precise definition of the statement f(x) is differentiable at x = c?

2. For which values of a ∈ R is

g(x) = \{ \begin{matrix}x^{a} & if x ∈ Q \\ 0 & if x ∉ Q \end{matrix}

differentiable at x = 0?